OpenStudy (anonymous):

Solve for q if 8/q = 64/16 4 2 8 16

OpenStudy (anonymous):

4 2 8 16

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

\[\frac{ 8 }{ q } = \frac{ 64 }{ 16 }\] cross multiply 64q = 128 q = 2

OpenStudy (unklerhaukus):

8 / q = 64 / 16 start by simplifying the right hand side 8 / q = 4 x 16 / 16 8 / q = 4 now multiply both sides by q q x 8 / q = 4q 8 = 4q and finally divid both sides by 4 8 / 4 = 4q / 4 2 x 4 / 4 = q 2 =q